OpenStudy (anonymous):

A hockey season ticket holder pays $72.48 total for her tickets plus $6.00 for a program each time she comes to a game, while another fan pay $18.08 for each ticket but doesn’t have to pay for their programs. a) Identify a variable for this equation. What does the variable represent? b) Create an equation for this situation to solve for the variable c) Solve the equation, find the value of the variable, check your work

OpenStudy (kropot72):

Notice that the season ticket holder buys "tickets" (plural). What do you think the variable is?

OpenStudy (anonymous):

Wat do u mean Variable...? is it the amount of tickets?

OpenStudy (kropot72):

Yes it is the number of tickets for the season of games. We can let the variable = n.

OpenStudy (anonymous):

oh ok, than a) is Variable = n and n = something right ?

OpenStudy (anonymous):

and n is anything we want right ?

OpenStudy (kropot72):

We have to find a value for n from the given information. Lets assume that the cost of a ticket works out to be the same for the season of tickets as for a single ticket. Then 72.48/n = 18.08 * n

OpenStudy (kropot72):

Multiplying both sides of the equation by n gives: 72.48 = 18.08 * n^2 n = square root of (72.48/18.08) Can you use a calculator to find n ?

OpenStudy (anonymous):

...

OpenStudy (anonymous):

wow i's so complex

OpenStudy (kropot72):

Not really. 72.48/18.08 = 4. The square root of 4 is 2. So the season is just 2 games.

OpenStudy (anonymous):

wai is it n^2 though ? can it be like 72.48n=18.08n

OpenStudy (kropot72):

72.48/n = 18.08 * n When 72.48/n is multiplied by n you get 72.48 When 18.08 * n is multiplied by n you get 18.08 * n^2 72.48 = 18.08 * n^2 Do you understand each step?

OpenStudy (anonymous):

yep

OpenStudy (kropot72):

Good :).

OpenStudy (anonymous):

so for b) 72.48/n = 18.08 * n <-- that is the equation rite ?

OpenStudy (anonymous):

72.48/18.08 = 4. The square root of 4 is 2. So the season is just 2 games. and that is the solution rite ?

OpenStudy (kropot72):

Yes and Yes.